Circular No. 7284

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1859+226 S. Chaty, C. A. Haswell, and A. J. Norton, Open University; R. I.

Hynes, Southampton University; I. Smail, Durham University; J.-E. Solheim and R. Ostensen, Tromsø University; T. R. Greve and J. Fynbo, Copenhagen University; K. Horne, K. O’Brien, and W. Skidmore, University of St. Andrews; R. Fried, Flagstaff, AZ; K. Krisciunas, University of Washington; P. M. Garnavich, M. Garcia, and M. Contreras, Harvard-Smithsonian Center for Astrophysics; E. T. Harlaftis, National Observatory, Athens; P. A. Charles and T. Shahbaz, Oxford; E. Kuulkers, Space Research Organization Netherlands; W. Chen and C. R. Shrader, Goddard Space Flight Center, NASA; S. B. Howell, University of Wyoming; and M. Wagner, Ohio State University, report the following near-infrared photometry and follow-up optical photometry of the counterpart to the x-ray transient XTE J1859+226 (IAUC 7274, 7279): Oct. 14.2 UT, J = 14.508 ± 0.011, H = 14.231 ± 0.014, K = 13.833 ± 0.013 (3.8-m U.K. Infrared Telescope + UFTI). Oct. 13.90, B = 15.9 ± 0.1; Oct. 14.85, B = 15.407 ± 0.034, V = 15.478 ± 0.035, R = 15.13 ± 0.11, I = 14.904 ± 0.027; Oct. 16.9, B = 16.34 ± 0.17, V = 15.603 ± 0.043, R = 15.324 ± 0.042, I = 15.005 ± 0.050 (2.5-m Nordic Optical Telescope at La Palma). Oct. 14.11, B = 16.134, V = 15.75, R = 15.294 (Braeside Observatory 0.40-m telescope at Flagstaff; uncertainties < 0.03). Oct. 14.1173, V = 15.501 ± 0.013, B − V = +0.344 ± 0.023; Oct. 14.1340, V = 15.484 ± 0.013, B − V = +0.388 ± 0.023; Oct. 15.1178, V = 15.329 ± 0.010, B − V = +0.377 ± 0.017, V −R = +0.267 ± 0.016, V −I = +0.537 ± 0.019 (Manashtash Ridge Observatory 0.76-m telescope). Oct. 15.1, R (Kitt Peak Harris) = 15.06 ± 0.02 (Fred L. Whipple Observatory 1.2-m telescope). Observations with the 0.94-m James Gregory Telescope at St. Andrews on Oct. 13.9 showed that, at V = 15.4 ± 0.15, the source exhibited evidence of modulation through a 0.04-mag amplitude.

SUPERNOVA 1999ee IN IC 5179 A. Williams, Perth Observatory, reports that J. Biggs and A. Verveer

obtained four unfiltered CCD images of SN 1999ee (cf. IAUC 7272) around Oct. 18.49 UT with the Perth 0.25-m Mike Candy Telescope, yielding the following precise position for the supernova: α = 22^{h}16^{m}10^{s}.00 ± 0^{s}.02, δ = −36^{o}50′39′′.7 ± 0′′.2 (equinox 2000.0; GSC reference stars), which is ∼ 13′′ east and 3′′.5 south of the galaxy’s center; the magnitude was estimated as 15.5 ± 0.5.

1999 October 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green